OpenStudy (anonymous):

Simplify by factoring. Assume that all expressions under radicals represent nonnegative numbers √75a^2b

OpenStudy (anonymous):

I get 75ab

OpenStudy (anonymous):

\[75=25\times 3\] and \[\sqrt{a^2}=a\]so you get \[5a\sqrt{3b}\]

OpenStudy (anonymous):

slowly \[\sqrt{75ab^2}=\sqrt{25\times b^2\times 3\times a}=\sqrt{25}\sqrt{b^2}\sqrt{3a}=5b\sqrt{3a}\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw